OpenStudy (anonymous):

What is the equation of the line passing through (–3, –9) and parallel to the line y = x + 6 in standard form?

OpenStudy (shubhamsrg):

two parallel lines have the same slope slope of y= x+6 is 1 as in y=mx + c,,m is the slope now the required eqn is y= x + c for some constant c since this line passes through (-3,-9) ,,substitute x and y for these values and find c..thus you should now have your ans..